Find the number c that satisfies the conclusion of th f(x) = e-2x [0,3] Submit Answer Practice Another Version
Answer #1

f(2)= fCC) = 1 - 2e-20 e 22 [0,37 f(b)-fa). . b-a . e-6-1 -6e-2C = 0-2 = -20 = e-6-1 e-6-1 = 1-go ln (1-60) 108 + entiende [